Determine the power dissipated when a current of 10Ma flows through an appliance having a resistance of 8kΩ
1
Expert's answer
2020-05-04T12:49:53-0400

"I=10^7A"

"R=8\times10^3 Ohm"

Power dissipated is "P=I^2R=8\times10^{17} W"